Let P(A) = 0.6, P(B) = 0.3 and P(A / B) = 0.5. Find P(A or B).
Answer by ikleyn(52776)

.
P(A | B)  actually means and is the same as  P(A & B)/P(B),  where &  represents the logical  "AND", or the intersection of the two sets.


Therefore, the given  P(A | B) = 0.5  means that  P(A & B)/P(B) = 0.5.


Multiplying both sides of the last equality by P(B), you get


    P(A & B) = 0.5*P(B) = 0.5*0.3 = 0.15.


Now last step is to use the general formula


    P(A or B) = P(A) + P(B) - P(A & B) = 0.6 + 0.3 - 0.15 = 0.75.    ANSWER

Solved.